two trains one of length 100m and another of length 125m are moving in mutually opposite directions along parallel lines, meet each other, each with speed 10m/s. if their acceleration are 0.3 m/s2 and 0.2m/s2 respectively then find the time taken to pass each other.

Grade:11

2 Answers

Deepak Kumar
6 years ago
Length of first train = 100m.
Length of second train = 125m.
So, total crossing distance = 100+125 = 225m.
Velocity of both train = 10m/s .
acceleration of first train = 0.3m/s2.
acceleration of second train = 0.2m/s2.
Since both train moving to opposite direction of each other.
 so, velocity and acceleration is added to find the time taken to cross each other.
so, velocity = 10+10 = 20m/s.
      acceleration = 0.3+0.2 = 0.5m/s.
Now using kinemaics second equation.
 S = Ut +at2/2.
225 = 20t +0.5t2/2.
  t+80t – 1000= 0.
Now solving aove equation.
  t = 10.99s and -90.99s.
negative value is not possible so t = 10.99 s.
 
Ajinkya
5 years ago
Do the same as he has done but in the place of 1000 there will be 900.
t^2+80t-900=0
t^2+90t-10t-900=0
t(t+90)-10(t+90)=0
t=-90 and t=10.
Here t=-90 is rejected so t=10 is the answer.